OpenStudy (anonymous):

two planes left an airport at noon, one flew east and the other flew west at twice the speed. After 3 hours the planes were 2700mi apart. How fast was each plane flying?

OpenStudy (michele_laino):

I call with v_A and v_B the velocities of each plane say plane A and plane B, furthermore, be: v_B=2*v_A then we can write: \[\Large d = {v_A}t + {v_B}t = {v_A}t + 2{v_A}t = 3{v_A}t\] where t is the time namely t= 3 hours, and d is their separation, namely d= 2,700 mi 

OpenStudy (michele_laino):

so, substituting your data we get: \[\Large 2700 = 3{v_A} \times 3\] please solve for v_A

OpenStudy (anonymous):

wait what?

OpenStudy (anonymous):

can u explain please

OpenStudy (michele_laino):

we have: \[\Large 2700 = 9{v_A}\]

OpenStudy (michele_laino):

now, please divide both sides by 9, what do you get?

OpenStudy (michele_laino):

hint:  what is v_A?